OpenStudy (anonymous):

slope 3/4, x-intercept -2 point-slope form

OpenStudy (anonymous):

You need to plot this on a graph. Then get two point, after that you can insert into the correct slots.

OpenStudy (anonymous):

how do I plot? Im confused

OpenStudy (anonymous):

Do you have graph paper?

OpenStudy (anonymous):

Isn't there a formula where I just plug in the numbers?

OpenStudy (anonymous):

First you need two points.

OpenStudy (anonymous):

so what exactly do i graph?

OpenStudy (anonymous):

Graph x-intercept 2. (2,0)

OpenStudy (anonymous):

okay now what?

OpenStudy (anonymous):

Now use the slope formula. every time you go up 3, go to the right 4.

OpenStudy (anonymous):

now what? i did that

OpenStudy (anonymous):

Get two points from the slope.

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

Then plug into y - y1 = m(x - x1)

OpenStudy (anonymous):

Where m is the slope.

OpenStudy (anonymous):

x1 = -2 y1 = 0

OpenStudy (anonymous):

so plug in the two points into that formula? now what

OpenStudy (anonymous):

y - 0 = 3/4(x + 2)

OpenStudy (anonymous):

That's it.

OpenStudy (ciel21):

its purple :3